# Field-sensitive dislocation bound states in two-dimensional $d$-wave altermagnets

### v1

@@ Abstract
When a two-dimensional $d$-wave altermagnet is grown on a substrate, the interplay of momentum-dependent spin splittings arising from altermagnetism and Rashba spin-orbit coupling gives rise to a nodal band structure with band degeneracies enforced by a $C_{4z}\mathcal{T}$ symmetry. If we break the $C_{4z}\mathcal{T}$ symmetry by an exchange field, the band degeneracies are found to be immediately lifted, leading to a topological band structure characterized by nontrivial strong and weak topological indices. Remarkably, both the strong topological index and the $Z_{2}$-valued weak topological indices depend sensitively on the direction of the exchange field. As a consequence of the bulk-defect correspondence, we find that the unique dependence of weak topological indices on the exchange field in this system dictates that the presence or absence of topological bound states at lattice dislocations also depends sensitively on the direction of the exchange field. When the substrate is an $s$-wave superconductor, we find that a similar dependence of band topology on the exchange field gives rise to field-sensitive dislocation Majorana zero modes. As topological dislocation bound states are easily detectable by scanning tunneling microscopy, our findings unveil a promising experimental diagnosis of altermagnetic materials among an ever growing list of candidates.
@@ I Introduction
Altermagnetism (AM) has attracted increasing interest recently as a collinear magnetic order with salient properties distinct from the conventional collinear ferromagnetism and antiferromagnetismŠmejkal et al. (2022a, b, 2020); Hayami et al. (2019, 2020); Yuan et al. (2020, 2021); Mazin et al. (2021); Shao et al. (2021); Ma et al. (2021); Liu et al. (2022); Feng et al. (2022); Gonzalez Betancourt et al. (2023); Mazin (2023); Turek (2022); Hariki et al. (2023); Bhowal and Spaldin (2024); Zhou et al. (2024); Han et al. (2024); Chen et al. (2023); Xiao et al. (2023); Jiang et al. (2023). In real space, the magnetic moments of an altermagnet are collinearly arranged to form a Néel order, just like an antiferromagnet. However, unlike the antiferromagnets the two sublattices with opposite magnetic moments in a magnetic unit cell of the altermagnet cannot be mapped to each other by the combined symmetry operations of time reversal and inversion/translation. Instead, they are mapped to each other by the combined symmetry operations of time reversal and rotation/mirrorŠmejkal et al. (2022a, b). A remarkable consequence of this difference in symmetry is that the AM leads to momentum-dependent spin-splitting electronic band structures but maintains symmetry-compensated zero net magnetization. Thus, it is not only distinct from ferromagnetism which leads to momentum-independent spin splitting and finite magnetization, but also distinguishes itself from antiferromagnetism which results in degenerate band structures. Notably, the spin splitting induced by AM can reach the order of $1$eV, and the symmetry pattern of the spin splitting is rich, exemplified by the classification of AM into groups dubbed $d$-, $g$- and $i$-wave AM. Excitingly, several materials, including insulating MnTeOsumi et al. (2024); Lee et al. (2024); Krempaský et al. (2024); Hajlaoui et al. (2024) and metallic CrSbReimers et al. (2024); Ding et al. (2024); Yang et al. (2024); Zeng et al. (2024); Li et al. (2024), have been experimentally confirmed to be altermagnets by high-resolution angle-resolved photoemission spectroscopy (ARPES). On the theoretical front, first-principle Guo et al. (2023); Gao et al. (2023); Qu et al. (2024); Sødequist and Olsen (2024); Bernardini et al. (2024); Sheoran and Bhattacharya (2024); Milivojević et al. (2024); Zeng and Zhao (2024); Liu et al. (2024) and model Leeb et al. (2024); Das et al. (2023); Sato et al. (2023); Roig et al. (2024); Cônsoli and Vojta (2024); Yu et al. (2024); Bose et al. (2024) calculations have predicted an ever growing list of candidates for AM. Furthermore, many studies have also shown that AM can give rise to numerous interesting effects and phases, such as giant and tunneling magnetoresistanceŠmejkal et al. (2022c), diverse tunneling phenomena in superconductor/altermagnet junctionsOuassou et al. (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Cheng and Sun (2024); Nagae et al. (2024); Gløckner Giil et al. (2024); Das and Soori (2024); Lu et al. (2024), finite-momentum Cooper pairingZhang et al. (2024); Sumita et al. (2023); Chakraborty and Black-Schaffer (2023), unconventional superconductivityZhu et al. (2023); Wei et al. (2024); Brekke et al. (2023); Mæland et al. (2024); Chourasia et al. (2024), various types of Hall effectsParshukov et al. (2024); Fang et al. (2023); Attias et al. (2024); Hoyer et al. (2024), and anisotropic RKKY interaction between spin impuritiesLee (2023); Amundsen et al. (2024).
@@
Because altermagnetic materials are often grown on substrates which can induce spin-orbit coupling (SOC), theoretical studies of altermagnetic materials need to also take into account this effect. SOC is known as another basic mechanism giving rise to momentum-dependent spin splitting and the interplay of AM and SOC can result in band structures of rich topological propertiesFernandes et al. (2024). Thus far, the $d$-wave AM has been the focus of most of the theoretical studies. It has been shown that the band structure of an intrinsic $d$-wave altermagnet normally has spin-polarized Dirac points in two dimensions (2D), and the further presence of SOC can gap out the Dirac points and result in first-order topological insulator phasesAntonenko et al. (2024); Parshukov et al. (2024). Furthermore, when intrinsic superconductivity occurs in a 2D $d$-wave altermagnet with Rashba SOC, both first-order and second-order topological superconductivity are found to emergeZhu et al. (2023). Lastly, it has been shown that second-order topological insulators or superconductors can be obtained in hybrid systems composed of $d$-wave altermagnets and first-order topological insulators or superconductorsGhorashi et al. (2023); Li and Liu (2023); Li et al. (2024); Li (2024); Ezawa (2024). In all these studies, the nontrivial momentum-space topology of the bulk bands is manifested in the presence of topological boundary states, dictated by the bulk-boundary correspondence. Interestingly, real-space topological defects, which are ubiquitous in materials and are characterized by real-space topological invariants, can also reflect the momentum-space topology in a distinct way, known as the bulk-defect correspondenceTeo and Kane (2010); Chiu et al. (2016); Yao et al. (2017).
@@
The types of topological defects in materials are diverseMermin (1979); Kleman and Friedel (2008). For lattice topological defects, disclinations and dislocations are two classes that attract particular interest in the context of topological phasesTeo and Hughes (2017); Lin et al. (2023); Asahi and Nagaosa (2012); Bühler et al. (2014); Hughes et al. (2014); Slager et al. (2014); de Juan et al. (2014); van Miert and Ortix (2018); Roy and Juričić (2021); Soto-Garrido et al. (2020); Schindler et al. (2022); Hu and Zhang (2024). In the seminal work of Ran, Zhang and VishwanathRan et al. (2009), it was discovered that a lattice dislocation can harbor a pair of 1D gapless helical modes in 3D topological insulators; the topological criterion for the existence of these topological dislocation modes is given by $\textbf{B}\cdot\textbf{M}_{\nu}=\pi\,(\text{mod}\,2\pi)$, where B refers to the Burgers vector which is the real-space topological invariant characterizing the dislocation, and $\textbf{M}_{\nu}=\sum_{i=1}^{3}\nu_{i}\textbf{G}_{i}/2$. Here $\nu_{i}=\{0,1\}$ is the weak topological indices defined on the high symmetry planes at the Brillouin zone’s boundaryFu and Kane (2007) and $\textbf{G}_{i}$ is the reciprocal lattice vectors. $\textbf{M}_{\nu}$ acts like a time-reversal invariant momentum. Later, it was demonstrated that the topological criterion can also be generalized to 2D and is applicable to topological superconductors as wellRan (2010). In 2D, when the topological criterion is fulfilled, the topological dislocation modes are 0D bound states. As an application, it was shown that the presence or absence of dislocation bound states in a 2D topological insulator with $C_{4z}$ rotation symmetry can serve as a bulk probe to diagnose the location of band inversionJuričić et al. (2012).
@@
Figure 1: (a) A schematic diagram of field-sensitive dislocation bound states and their experimental measurement by scanning tunneling microscopy. The 2D altermagnetic material grown on an insulator substrate or an $s$-wave superconductor has two pairs of dislocations with perpendicular Burgers vectors. The color correspondence between the bell-shaped surfaces denoting topological dislocation bound states and the arrows denoting the exchange field’s direction illustrates the fact that which pair of the dislocations carries topological bound states relies on the exchange field’s direction. (b) Details of the dislocation configuration. The two blue (red) dots correspond to the cores of a pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ (±$\pm\bm{e_{y}}$). The two pairs of dislocations with perpendicular Burgers vectors form a cross-shaped defect. (c) 2D Brillouin zone. $\boldsymbol{\Gamma}=(0,0)$, $\textbf{X}=(\pi,0)$, $\textbf{Y}=(0,\pi)$ and $\textbf{M}=(\pi,\pi)$ are four high symmetry points. Two weak topological indices, $\nu_{x}$ and $\nu_{y}$, which determine the presence or absence of topological dislocation bound states, are defined on the blue and red lines at the Brillouin zone’s boundary.
@@ II Dislocation bound states in 2D metal with $d$-wave AM and Rashba SOC
Here $\sigma_{i}$ are Pauli matrices acting on the spin degrees of freedom; the first term in $\mathcal{H}_{0}$ is the kinetic energy arising from the nearest-neighbor hoppings, the second term denotes the Rashba SOC originating from structure asymmetry, and the last term accounts for the presence of two types of exchange fields, where the first momentum-dependent part ($t_{\rm AM}$) is attributed to $d$-wave AM and the second momentum-independent part ($M_{z}$) to an external perpendicular magnetic field or a ferromagnetic insulator substrate whose magnetization is perpendicular to the plane. Throughout the lattice constant is set to unity for notational simplicity, and without loss of generality the parameters $t$, $\lambda$ and $t_{\rm AM}$ are assumed to be non-negative for the convenience of discussion.
@@
$C_{\pm}$ $=$ ±$\pm\frac{1}{2\pi}\int_{BZ}\frac{\bm{d}(\bm{k})\cdot[\partial_{k_{x}}\bm{d}(\bm{k})\times\partial_{k_{y}}\bm{d}(\bm{k})]}{2|\bm{d}(\bm{k})|^{3}}d^{2}k$ (2)
$=$ ±$\pm\begin{cases}{\rm sgn}(M_{z}),&0<|M_{z}|<4t_{\rm AM},\\ 0,&|M_{z}|>4t_{\rm AM}.\end{cases}$
where the subscript +/+/-+ / - refers to the upper/lower band, and $\bm{d}(\boldsymbol{k})=(2\lambda\sin k_{y},-2\lambda\sin k_{x},2t_{\rm AM}(\cos{k_{x}}-\cos{k_{y}})+M_{z})$, with the components $d_{i}(\boldsymbol{k})$ corresponding to the coefficient functions in front of the Pauli matrix $\sigma_{i}$ in Eq. (1). The dependence of Chern number on $M_{z}$ suggests that an arbitrarily weak perpendicular exchange field render the band structure topologically nontrivial. Furthermore, the Chern number has a sensitive dependence on the direction of the exchange field. Thus, a reversal of the direction of the exchange field will change the sign of the Chern number. As topological phases have bulk-boundary correspondence, it is natural to expect that the topological boundary states will also have an intriguing dependence on the exchange field.
By calculating the energy spectrum for a sample of ribbon geometry, we find as expected that the chiral edge state reverses its chirality when the exchange field reverses its direction. What is surprising, however, is that the momentum at which the edge-state spectrum crosses undergoes a jump. When $M_{z}>0$, the edge-state spectrum crossing occurs at $k_{x}=\pi$ on the $y$-normal edges [Fig.2(a)] and at $k_{y}=0$ on the $x$-normal edges [Fig.2(b)]. Reversing the direction of the exchange field, the crossing momentum has a jump of half the reciprocal lattice vector. That is, the edge-state spectrum crossing is shifted to $k_{x}=0$ on the $y$-normal edges [Fig.2(c)] and to $k_{y}=\pi$ on the $x$-normal edges [Fig.2(d)]. The edge-state spectrum crossing is connected to weak topological indices defined on one-dimensional lower submanifolds of the Brillouin zone, namely, the high symmetry lines of the Brillouin zone. The jump of the crossing momentum then suggests that the weak topological indices of this Hamiltonian also have a sensitive dependence on the direction of the exchange field.
@@
where $\xi(\bm{k}_{R})$ is the $C_{2z}$ operator’s eigenvalue for the lower-band eigenstate at the $C_{2z}$-invariant momentum $\bm{k}_{R}\in\{\textbf{X},\textbf{Y},\textbf{M}\}$, i.e., $C_{2z}|u(\bm{k}_{R})\rangle=\xi(\bm{k}_{R})|u(\bm{k}_{R})\rangle$. Because these eigenvalues take values of ±$\pm i$, a factor of “$-1$” is introduced on the right hand side of the above equations. A straightforward calculation reveals
@@
Figure 3: Phase diagram determined via weak topological indices. The red and blue regions correspond to $(\nu_{x},\nu_{y})=(0,1)$ and $(1,0)$, respectively. The two insets below the phase diagram are energy spectra for a system with periodic boundary conditions in both $x$ and $y$ directions. The results show that, in the region with $(\nu_{x},\nu_{y})=(0,1)$, dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ harbor bound states, while the ones with $\textbf{B}=\pm\bm{e_{x}}$ do not, and the situation is just the opposite for the region with $(\nu_{x},\nu_{y})=(1,0)$. For the left (right) inset, $t_{\rm AM}=0.3$ and $M_{z}=-0.24$ $(0.24)$. The values of other parameters are $t=0$ and $\lambda=0.3$.
@@
To verify this expectation, we place a cross-shaped defect consisting of two pairs of dislocations in the system [see Fig.1(b)] and numerically diagonalize the Hamiltonian under periodic boundary conditions in both $x$ and $y$ directions. In the case of $0<M_{z}<4t_{\rm AM}$ for which $(\nu_{x},\nu_{y})=(1,0)$, we find that the pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ harbor topological bound states while those with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ do not, as shown in Fig.3. The picture is just reversed when $-4t_{\rm AM}<M_{z}<0$, demonstrating that the presence or the absence of topological bound states at a dislocation can be tuned by simply adjusting the exchange field’s direction.

### v2

@@ Abstract
When a two-dimensional $d$-wave altermagnet is grown on a substrate, the interplay of momentum-dependent spin splittings arising from altermagnetism and Rashba spin-orbit coupling gives rise to a nodal band structure with band degeneracies enforced by a $C_{4z}\mathcal{T}$ symmetry. If the $C_{4z}\mathcal{T}$ symmetry is broken by an exchange field, we find that the band degeneracies are immediately lifted, leading to a topological band structure characterized by nontrivial strong and weak topological indices. Remarkably, both the strong topological index and the $Z_{2}$-valued weak topological indices depend sensitively on the direction of the exchange field. As a consequence of the bulk-defect correspondence, we find that the unique dependence of weak topological indices on the exchange field in this system dictates that the presence or absence of topological bound states at lattice dislocations also depends sensitively on the direction of the exchange field. When the substrate is an $s$-wave superconductor, we find that a similar dependence of band topology on the exchange field gives rise to field-sensitive dislocation Majorana zero modes. As topological dislocation bound states are easily detectable by scanning tunneling microscopy, our findings unveil a promising experimental diagnosis of altermagnetic materials among an ever growing list of candidates.
@@ I Introduction
Altermagnetism (AM) has attracted increasing interest recently as a collinear magnetic order with salient properties distinct from the conventional collinear ferromagnetism and antiferromagnetismŠmejkal et al. (2022a, b, 2020); Hayami et al. (2019, 2020); Yuan et al. (2020, 2021); Mazin et al. (2021); Shao et al. (2021); Ma et al. (2021); Liu et al. (2022); Feng et al. (2022); Gonzalez Betancourt et al. (2023); Mazin (2023); Turek (2022); Hariki et al. (2023); Bhowal and Spaldin (2024); Zhou et al. (2024); Han et al. (2024); Chen et al. (2023); Xiao et al. (2023); Jiang et al. (2023). In real space, the magnetic moments of an altermagnet are collinearly arranged to form a Néel order, just like an antiferromagnet. However, unlike the antiferromagnets the two sublattices with opposite magnetic moments in a magnetic unit cell of the altermagnet cannot be mapped to each other by the combined symmetry operations of time reversal and inversion/translation. Instead, they are mapped to each other by the combined symmetry operations of time reversal and rotation/mirrorŠmejkal et al. (2022a, b). A remarkable consequence of this difference in symmetry is that the AM leads to momentum-dependent spin-splitting electronic band structures but maintains symmetry-compensated zero net magnetization. Thus, it is not only distinct from ferromagnetism which leads to momentum-independent spin splitting and finite magnetization, but also distinguishes itself from antiferromagnetism which results in degenerate band structures. Notably, the spin splitting induced by AM can reach the order of 11eV, and the symmetry pattern of the spin splitting is rich, exemplified by the classification of AM into groups dubbed $d$-, $g$- and $i$-wave AM. Excitingly, several materials, including insulating MnTeOsumi et al. (2024); Lee et al. (2024); Krempaský et al. (2024); Hajlaoui et al. (2024) and metallic CrSbReimers et al. (2024); Ding et al. (2024); Yang et al. (2024); Zeng et al. (2024); Li et al. (2024), have been experimentally confirmed to be altermagnets by high-resolution angle-resolved photoemission spectroscopy (ARPES). On the theoretical front, first-principle Guo et al. (2023); Gao et al. (2023); Qu et al. (2024); Sødequist and Olsen (2024); Bernardini et al. (2024); Sheoran and Bhattacharya (2024); Milivojević et al. (2024); Zeng and Zhao (2024); Liu et al. (2024) and model Leeb et al. (2024); Das et al. (2023); Sato et al. (2023); Roig et al. (2024); Cônsoli and Vojta (2024); Yu et al. (2024); Bose et al. (2024) calculations have predicted an ever growing list of candidates for AM. Furthermore, many studies have also shown that AM can give rise to numerous interesting effects and phases, such as giant and tunneling magnetoresistanceŠmejkal et al. (2022c), diverse tunneling phenomena in superconductor/altermagnet junctionsOuassou et al. (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Cheng and Sun (2024); Nagae et al. (2024); Gløckner Giil et al. (2024); Das and Soori (2024); Lu et al. (2024), finite-momentum Cooper pairingZhang et al. (2024); Sumita et al. (2023); Chakraborty and Black-Schaffer (2023), unconventional superconductivityZhu et al. (2023); Wei et al. (2024); Brekke et al. (2023); Mæland et al. (2024); Chourasia et al. (2024), various types of Hall effectsParshukov et al. (2024); Fang et al. (2023); Attias et al. (2024); Hoyer et al. (2024), and anisotropic RKKY interaction between magnetic impuritiesLee (2023); Amundsen et al. (2024).
@@
Because altermagnetic materials are often grown on substrates which can induce spin-orbit coupling (SOC), theoretical studies of altermagnetic materials need to also take into account this effect. SOC is known as another basic mechanism giving rise to momentum-dependent spin splitting and the interplay of AM and SOC can result in band structures of rich topological propertiesFernandes et al. (2024). Thus far, the $d$-wave AM has been the focus of most of the theoretical studies. It has been shown that the band structure of an intrinsic $d$-wave altermagnet normally has spin-polarized Dirac points in two dimensions, and the further presence of SOC can gap out the Dirac points and result in first-order topological insulator phasesAntonenko et al. (2024); Parshukov et al. (2024). Furthermore, when intrinsic superconductivity occurs in a two-dimensional (2D) $d$-wave altermagnet with Rashba SOC, both first-order and second-order topological superconductivity are found to emergeZhu et al. (2023). Lastly, it has been shown that second-order topological insulators or superconductors can be obtained in hybrid systems composed of $d$-wave altermagnets and first-order topological insulators or superconductorsGhorashi et al. (2023); Li and Liu (2023); Li et al. (2024); Li (2024); Ezawa (2024). In all these studies, the nontrivial momentum-space topology of the bulk bands is manifested in the presence of topological boundary states, dictated by the bulk-boundary correspondence. Interestingly, real-space topological defects, which are ubiquitous in materials and are characterized by real-space topological invariants, can also reflect the momentum-space topology in a distinct way, known as the bulk-defect correspondenceTeo and Kane (2010); Chiu et al. (2016); Yao et al. (2017).
@@
The types of topological defects in materials are diverseMermin (1979); Kleman and Friedel (2008). For lattice topological defects, disclinations and dislocations are two classes that attract particular interest in the context of topological phasesTeo and Hughes (2017); Lin et al. (2023); Asahi and Nagaosa (2012); Bühler et al. (2014); Hughes et al. (2014); Slager et al. (2014); de Juan et al. (2014); van Miert and Ortix (2018); Roy and Juričić (2021); Soto-Garrido et al. (2020); Schindler et al. (2022); Hu and Zhang (2024). In the seminal work of Ran, Zhang and VishwanathRan et al. (2009), it was discovered that a lattice dislocation can harbor a pair of 1D gapless helical modes in 3D topological insulators; the topological criterion for the existence of these topological dislocation modes is given by $\textbf{B}\cdot\textbf{M}_{\nu}=\pi\,(\text{mod}\,2\pi)$, where B refers to the Burgers vector which is the real-space topological invariant characterizing the dislocation, and $\textbf{M}_{\nu}=\sum_{i=1}^{3}\nu_{i}\textbf{G}_{i}/2$. Here $\nu_{i}=\{0,1\}$ is the weak topological indices defined on the high symmetry planes at the Brillouin zone’s boundaryFu and Kane (2007) and $\textbf{G}_{i}$ is the reciprocal lattice vectors. $\textbf{M}_{\nu}$ acts like a time-reversal invariant momentum. Later, it was demonstrated that the topological criterion can also be generalized to two dimensions and is applicable to topological superconductors as wellRan (2010). In two dimensions, when the topological criterion is fulfilled, the topological dislocation modes are 0D bound states. As an application, it was shown that the presence or absence of dislocation bound states in a 2D topological insulator with $C_{4z}$ rotation symmetry can serve as a bulk probe to diagnose the location of band inversionJuričić et al. (2012).
@@
Figure 1: (a) A schematic diagram of field-sensitive dislocation bound states and their experimental measurement by scanning tunneling microscopy. The 2D altermagnetic material grown on an insulator substrate or an $s$-wave superconductor has two pairs of dislocations with perpendicular Burgers vectors. The color correspondence between the bell-shaped surfaces denoting topological dislocation bound states and the arrows denoting the exchange field’s direction illustrates the fact that which pair of the dislocations carries topological bound states relies on the exchange field’s direction. (b) Details of the dislocation configuration. The two blue (red) dots correspond to the cores of a pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ ($\pm\bm{e_{y}}$). The two pairs of dislocations with perpendicular Burgers vectors form a cross-shaped defect. (c) 2D Brillouin zone. $\boldsymbol{\Gamma}=(0,0)$, $\textbf{X}=(\pi,0)$, $\textbf{Y}=(0,\pi)$ and $\textbf{M}=(\pi,\pi)$ are four high symmetry points. Two weak topological indices, $\nu_{x}$ and $\nu_{y}$, which determine the presence or absence of topological dislocation bound states, are defined on the blue and red lines at the Brillouin zone’s boundary.
@@ II Dislocation bound states in 2D metal with $d$-wave AM and Rashba SOC
Here $\sigma_{i}$ are Pauli matrices acting on the spin degrees of freedom. This two-band Hamiltonian can emerge, for example, in the Lieb lattice. To be specific, when the sublattices at the centers of the $x$- and $y$-directional edges have opposite magnetic moments, the electrons from the nonmagnetic atoms at the vertices of the Lieb lattice will experience an effective $d$-wave altermagnetic exchange fieldBrekke et al. (2023). If one considers only the effective Hamiltonian for the nonmagnetic atoms, the first hopping term (the $t$ term) and the $d$-wave AM term (the $t_{\rm AM}$ term) in $\mathcal{H}_{0}$ will be obtained. The second term in $\mathcal{H}_{0}$ denotes the Rashba SOC originating from structure asymmetry, and the last $M_{z}$ term accounts for a momentum-independent exchange field induced by an external perpendicular magnetic field or a ferromagnetic insulator substrate whose magnetization is perpendicular to the plane. Throughout the lattice constant is set to unity for notational simplicity, and without loss of generality the parameters $t$, $\lambda$ and $t_{\rm AM}$ are assumed to be non-negative for the convenience of discussion.
@@
$C_{\pm}$ $=$ $\pm\frac{1}{2\pi}\int_{BZ}\frac{\bm{d}(\bm{k})\cdot[\partial_{k_{x}}\bm{d}(\bm{k})\times\partial_{k_{y}}\bm{d}(\bm{k})]}{2|\bm{d}(\bm{k})|^{3}}d^{2}k$ (2)
$=$ $\pm\begin{cases}{\rm sgn}(M_{z}),&0<|M_{z}|<4t_{\rm AM},\\ 0,&|M_{z}|>4t_{\rm AM}.\end{cases}$
where the subscript +/+/- refers to the upper/lower band, and $\bm{d}(\boldsymbol{k})=(2\lambda\sin k_{y},-2\lambda\sin k_{x},2t_{\rm AM}(\cos{k_{x}}-\cos{k_{y}})+M_{z})$, with the components $d_{i}(\boldsymbol{k})$ corresponding to the coefficient functions in front of the Pauli matrix $\sigma_{i}$ in Eq. (1). The dependence of Chern number on $M_{z}$ suggests that an arbitrarily weak perpendicular exchange field renders the band structure topologically nontrivial. Furthermore, the Chern number has a sensitive dependence on the direction of the exchange field. Thus, a reversal of the direction of the exchange field will change the sign of the Chern number. As topological phases have bulk-boundary correspondence, it is natural to expect that the topological boundary states will also have an intriguing dependence on the exchange field.
By considering a sample of ribbon geometry (periodic boundary conditions in $x$ direction and open boundary conditions in $y$ direction, or vise versa) and calculating the energy spectrum of the Hamiltonian in Eq.(1), we find as expected that the chiral edge state reverses its chirality when the exchange field reverses its direction. What is surprising, however, is that the momentum at which the edge-state spectrum crosses undergoes a jump. When $M_{z}>0$, the edge-state spectrum crossing occurs at $k_{x}=\pi$ on the $y$-normal edges [Fig.2(a)] and at $k_{y}=0$ on the $x$-normal edges [Fig.2(b)]. Reversing the direction of the exchange field, the crossing momentum has a jump of half the reciprocal lattice vector. That is, the edge-state spectrum crossing is shifted to $k_{x}=0$ on the $y$-normal edges [Fig.2(c)] and to $k_{y}=\pi$ on the $x$-normal edges [Fig.2(d)]. Interestingly, the crossing momenta on the $x$-normal edges and $y$-normal edge states are just switched upon reversing the exchange field. This switching behavior can be understood through the $C_{4z}\mathcal{T}$ symmetry. To be specific, for the $x$-normal edge states under positive $M_{z}$ as shown in Fig. 2(a), applying the $C_{4z}\mathcal{T}$ symmetry operation gives the $y$-normal edge states under negative $M_{z}$ as shown in Fig. 2(d), and vise versa. Similarly, the spectra in Fig. 2(b) and Fig. 2(c) are connected by the same symmetry operation. As the edge-state spectrum crossings at time-reversal invariant momenta are connected to weak topological indices defined on 1D lower submanifolds of the Brillouin zone, namely, high symmetry lines of the Brillouin zone, their dramatic jump suggests that the weak topological indices of this Hamiltonian also have a sensitive dependence on the direction of the exchange field.
@@
where $\xi(\bm{k}_{R})$ is the $C_{2z}$ operator’s eigenvalue for the lower-band eigenstate at the $C_{2z}$-invariant momentum $\bm{k}_{R}\in\{\textbf{X},\textbf{Y},\textbf{M}\}$, i.e., $C_{2z}|u(\bm{k}_{R})\rangle=\xi(\bm{k}_{R})|u(\bm{k}_{R})\rangle$. Because these eigenvalues take values of $\pm i$, a factor of “1-1” is introduced on the right hand side of the above equations. A straightforward calculation reveals
@@
Figure 3: Phase diagram determined via weak topological indices. The red and blue regions correspond to $(\nu_{x},\nu_{y})=(0,1)$ and $(1,0)$, respectively. The two insets in the middle are energy spectra for a system with lattice size $N_{x}\times N_{y}=29\times 29$ and periodic boundary conditions in both $x$ and $y$ directions. The two insets at the bottom show the corresponding distributions of the dislocation bound states in the lattice. The results show that, in the region with $(\nu_{x},\nu_{y})=(0,1)$, dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ harbor bound states, while the ones with $\textbf{B}=\pm\bm{e_{x}}$ do not, and the situation is just the opposite for the region with $(\nu_{x},\nu_{y})=(1,0)$. For the left (right) inset, $t_{\rm AM}=0.3$ and $M_{z}=-0.24$ $(0.24)$. The values of other parameters are $t=0$ and $\lambda=0.3$.
@@
To verify this expectation, we place a cross-shaped defect consisting of two pairs of dislocations in the system [see Fig.1(b)] and numerically diagonalize the Hamiltonian under periodic boundary conditions in both $x$ and $y$ directions. In the case of $0<M_{z}<4t_{\rm AM}$ for which $(\nu_{x},\nu_{y})=(1,0)$, we find that the pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ harbor topological bound states while those with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ do not, as shown in Fig.3. The picture is just reversed when $4t_{\rm AM}<M_{z}<0$, demonstrating that the presence or the absence of topological bound states at a dislocation can be tuned by simply adjusting the exchange field’s direction.